OpenStudy (anonymous):

Find the common ratio in a geometric sequence i which the first term is 1 and the seventh term is 4096.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[a_{n} = a_{1}r^{n - 1}\]\[a_{7} = a_{1}r^{7 - 1}\]\[4096 = 1(r^{6})\]\[4096 = r^{6}\]\[r = \sqrt[6]{4096}\]\[r = ?\]

OpenStudy (anonymous):

the sixth root of 4096 is 4